OpenStudy (chrisneedshelp):

.What is the center of the circle that has a diameter whose endpoints are (2, 7) and (-6, -1)? (4, 3) (-2, 4) (-2, 3)

OpenStudy (chrisneedshelp):

@d3Xter ?

hero (hero):

The center of the circle is the midpoint of the diameter segment.

OpenStudy (anonymous):

The question is asking you to find the mid points \((\dfrac{x_o + x_1}{2},\dfrac{ y_o+y_1}{2})\)

OpenStudy (anonymous):

\(\small\left(\dfrac{2-6}{2}, \dfrac{7-1}{2}\right)\) Mid point: \(\small-2, 3\)

OpenStudy (chrisneedshelp):

thanks guys

OpenStudy (anonymous):

You're welcome :)